OpenStudy (anonymous):

Solve 5x - 6y = -38 3x + 4y = 0

OpenStudy (anonymous):

a) (4, 3) b) (-4, 3) c) (4, -3) d) (-4, -3)

OpenStudy (jhannybean):

I hope you will actually attempt to do the problems today,because yesterday you were just spamming questions,looking for answers. You are not looking or trying to understand how people are solving the problems.

OpenStudy (anonymous):

Agreed^

OpenStudy (anonymous):

no im actually trying to do them when people help explain it.

OpenStudy (anonymous):

but the answers work to(;

jhonyy9 (jhonyy9):

b. look right

jhonyy9 (jhonyy9):

5x-6y=-38 multiplie first equation by 4 and second by 6 for elimine the term of y 3x+4y=0 20x-24y=-152 18x+24y=0 --------------- 38x 0 = -152 x= -152/38 = - 4 x = -4 5(-4)-6y=-38 -20-6y=-38 -6y=-18 y= -18/(-6) = 3 x= -4 y= 3